OpenStudy (anonymous):

Bobby has a bag containing 2 sizes of marbles. The big ones weigh 10 g and the small ones weigh 2 g. If there are five times as many small ones as big ones and the bag weighs 300 g, how many big marbles does he have?

OpenStudy (anonymous):

let small be s and big be b s=5b s(2)+b(10)=300 5b(2)+b(10)=300 10b+10b=300 20b=300 b=15 Big marbles are 15

OpenStudy (anonymous):

ok thank you I know you are busy but could you help me with one other type of problem?

OpenStudy (anonymous):

sure :)